The OSC is playing in self trigger mode, gets to the filter board, then C34-4.7u, next up R21-100k, after the R21, there is no signal anymore.
And now i just found out that IC2-TL072P was the wrong way around (which surprises me, I must have been tired when I put it in, I usually double check), it must have sourced alot of current because I noticed it got very very hot. burritoes hot! The power supply is still working, the AVR, display is, … and I also wondered why the display was so darkb4 ;). OMG
I wonder what other parts it fried, I changed the part (IC2) and still no output …
I try again tomorrow